Suppose a certain state university's college of business obtained the following results on the salaries of a recent graduating class:

Finance Majors Business Analytics Majors

n1 = 120

n2 = 30

x1 = $48,237

x2 = $55,417

s1 = $19,000

s2 = $10,000

(a)

Formulate hypotheses so that, if the null hypothesis is rejected, we can conclude that salaries for Finance majors are significantly lower than the salaries of Business Analytics majors. Use α = 0.05. (Let μ1 = the population mean salary for Finance majors, and let μ2 = the population mean salary for Business Analytics majors. Enter != for ≠ as needed.)

H0:

Ha:

(b)

What is the value of the test statistic? (Use

μ1 − μ2.

Round your answer to three decimal places.)

(c)

What is the p-value? (Round your answer to four decimal places.)

p-value =

(d)

What is your conclusion?

A. Reject H0. We can conclude that salaries for Finance majors are significantly lower than the salaries of Business Analytics majors.

B.Reject H0. We cannot conclude that salaries for Finance majors are significantly lower than the salaries of Business Analytics majors.   

C. Do not reject H0. We cannot conclude that salaries for Finance majors are significantly lower than the salaries of Business Analytics majors.

D.Do not reject H0. We can conclude that salaries for Finance majors are significantly lower than the salaries of Business Analytics majors.
Answer #1

a)

Ho:μ1>=μ2
Ha: μ12

b)

standard error se=√(S21/n1+S22/n2)= 2518.267
test stat t =(x1-x2o)/Se= -2.851

c)

p value : = 0.0027

(please try 0.0040 if this comes wrong)

A. Reject H0. We can conclude that salaries for Finance majors are significantly lower than the salaries of Business Analytics majors.